Find the equation of each line in the form of y=mx+b
The line through (3, 2) with an undefined slope?
How do I set this problem up?? Thank you for your help!

Answer by Alan3354(69443)

y-y1 = m*(x-x1) where (x1,y1) is the point
y-2 = m(x-3)
That's it.
------------
you can do:
y = mx-3m + 2, but that's no improvement.
Maybe y = m(x-3) + 2